OpenStudy (anonymous):

The sum of two consecutive odd integers is 52 . Find the integers

OpenStudy (nikita2):

x + (x+2) = 52 2x = 52 -2 x = 25 answ 25 27

OpenStudy (anonymous):

25 27?

OpenStudy (nikita2):

yes

OpenStudy (anonymous):

looks good to me